# Past, Present, and Future of Bug Tracking in the Generative AI Era

### v1

@@ 4.1.5. Evaluation
Evaluating AI Agents is inherently complex (Yehudai et al., 2025). Unlike traditional systems, where well-defined metrics such as accuracy or precision can be applied against a fixed gold standard, an agent-based workflow introduces fluidity that resists simple benchmarking. Many of the tasks performed by the agents lack a clear reference output, making it difficult to establish what constitutes a “correct” answer (Kapoor et al., 2024).
@@
This problem has been widely recognized in the literature on agent evaluation. For instance, benchmarks for AI agents are often fragile, non-standardized, and overfitted to narrow test cases, which prevents them from reflecting real-world utility (Kapoor et al., 2024). Similarly, surveys of LLM-based agents in software engineering stress that many agent tasks are inherently open-ended and multi-modal, making them difficult to evaluate using traditional gold-standard datasets (Wang et al., 2025e). The AgentBench (Liu et al., 2023) framework echoes this by highlighting that evaluating agents requires interactive, multi-turn, open-ended environments rather than static correctness checks.
@@ 4.1.7. Privacy Challenges of LLM-Powered Agents
One challenge lies in memory extraction, where adversaries can design prompts to force agents to reveal stored or historical user information. Studies have shown that private memory contents in agent systems can be exfiltrated through black-box interactions, raising concerns about retaining sensitive bug data (Wang et al., 2025b).
@@
Finally, the general opacity of agent behavior makes privacy risks harder to detect. Since agents continuously adapt their responses based on memory and prompts, unintentional disclosure of private bug reports or logs may go unnoticed until after data has been leaked. This unpredictability means that privacy risks are not isolated to adversarial cases but may also arise during normal operations (Zhang et al., 2025).
@@ 4.1.8. Economic Implications and Hidden Costs
Beyond the immediate technical and ethical challenges, the adoption of AI-powered bug tracking systems presents a complex set of economic considerations that go far beyond simple licensing fees. While initial cost-benefit analyses might focus on the efficiency gains from automation, they often fail to account for the substantial hidden costs. These include the financial overhead of API calls, which can escalate unexpectedly with increased usage, as well as the significant infrastructure and computational costs for organizations choosing to host and fine-tune models internally. The cost of integrating these new systems with existing legacy tools and workflows can also be a major expense, requiring extensive development and testing.
@@ 4.2.1. Adapting to Diverse Project Structures
Our proposed bug-tracking system assumes a certain team structure with dedicated roles for PMs, developers, and other team members. However, we acknowledge that this may not align with every project. The organizational structure of software development teams can vary significantly across different projects and companies. It is important for practitioners to adapt our system to fit their specific needs.
@@ 4.2.3. Modular System Architecture and Process Optimization
The components of our bug-tracking system are designed with a modular architecture, meaning each part, from bug report enhancement and reproduction to fault localization and patch generation, can be run either independently or as part of a single, unified workflow. This is a critical implication for practitioners, as it allows for significant flexibility and process optimization based on a project’s unique needs and constraints.
@@ 4.2.4. System Integration and Toolchain Adapters
Successfully integrating such a bug-tracking toolchain into existing systems requires a strategic approach focused on building flexible adapters and connectors. These components serve as a crucial layer, allowing the system to communicate with and leverage existing development tools without requiring a complete overhaul of an organization’s infrastructure. It is essential to develop these connectors for critical parts of the toolchain, such as VCSs like Git, which would allow the system to pull code and submit AI-generated patches as pull requests. Similarly, adapters for integrated development environments like Visual Studio Code [^28] https://code.visualstudio.com/ or IntelliJ [^29] https://www.jetbrains.com/idea/ would provide developers with real-time AI recommendations and debugging insights directly within their coding environment. Furthermore, building connectors for CI/CD platforms such as Jenkins [^30] https://www.jenkins.io/ or GitHub Actions is vital for automating the testing and validation of AI-generated fixes, ensuring that only verified code is merged. Finally, adapters for pre-existing bug-tracking systems like Jira would allow for the seamless import of bug reports, enabling a phased and non-disruptive adoption. This approach ensures that the new system acts as a powerful enhancement to the existing ecosystem, maximizing the return on the integration effort.
@@ 4.3.3. Agent Capabilities as Research Problems
As illustrated in Figure 2, each capability represented in the workflow, such as bug report enhancement, reproduction, classification, localization, patch generation, and verification, can itself be viewed as a distinct research problem. Advances in natural language processing can refine bug report enhancement, program analysis techniques can strengthen localization, and test generation research can improve reproduction and verification. Researchers can therefore explore each of the boxes in Figure 2 as independent research topics, while also examining how these functions interoperate within an end-to-end framework that spans the entire bug tracking system.
@@ 4.3.4. Domain-Specific Challenges
The effectiveness of automated bug tracking will vary across domains. In delivery systems, bugs often occur immediately, making them straightforward to detect and reproduce. In contrast, interactive environments such as gaming may require long execution traces or rare conditions before a bug surfaces, complicating both reproduction and verification. These contrasts suggest that domain-specific adaptations of automation strategies are a promising research direction.

### v2

@@ 4.1.6. Agentic Evaluation Challenges
Evaluating LLM/AI Agents is inherently complex (Yehudai et al., 2025). Unlike traditional systems, where well-defined metrics such as accuracy or precision can be applied against a fixed gold standard, an agent-based workflow introduces fluidity that resists simple benchmarking. Many of the tasks performed by the agents lack a clear reference output, making it difficult to establish what constitutes a “correct” answer (Kapoor et al., 2024).
@@
This problem has been widely recognized in the literature on agent evaluation. For instance, benchmarks for LLM/AI agents are often fragile, non-standardized, and overfitted to narrow test cases, which prevents them from reflecting real-world utility (Kapoor et al., 2024). Similarly, surveys of LLM-based agents in software engineering stress that many agent tasks are inherently open-ended and multi-modal, making them difficult to evaluate using traditional gold-standard datasets (Wang et al., 2025e). The AgentBench (Liu et al., 2023) framework echoes this by highlighting that evaluating agents requires interactive, multi-turn, open-ended environments rather than static correctness checks.
@@ 4.1.8. Privacy Challenges of LLM-Powered Agents
One challenge lies in memory extraction, where adversaries can design prompts to force agents to reveal stored or historical user information. Studies have shown that private memory contents in agent systems can be exfiltrated through black-box interactions, raising concerns about retaining sensitive bug data (Wang et al., 2025a).
@@
Finally, the general opacity of agent behavior makes privacy risks harder to detect. Since agents continuously adapt their responses based on memory and prompts, unintentional disclosure of private bug reports or logs may go unnoticed until after data has been leaked. This unpredictability means that privacy risks are not isolated to adversarial cases but may also arise during normal operations (Zhang et al., 2025a). To reduce these risks, future AI-powered bug tracking frameworks must carefully control what agents are allowed to store, access, and expose by enforcing strict data-scoping policies, isolating sensitive information across agent components, and incorporating continuous auditing mechanisms that monitor agent interactions for unintended disclosure before privacy breaches propagate.
@@ 4.1.9. Economic Implications and Hidden Costs
Beyond the immediate technical and ethical challenges, the adoption of AI-powered bug tracking systems presents a complex set of economic considerations that go far beyond simple licensing fees. While initial cost-benefit analyses might focus on the efficiency gains from automation, they often fail to account for the substantial hidden costs. These include the financial overhead of API calls, which can escalate unexpectedly with increased usage, as well as the significant infrastructure and computational costs for organizations choosing to host and fine-tune models internally. The cost of integrating these new systems with existing legacy tools and workflows can also be a major expense, requiring extensive development and testing. To mitigate these economic risks, organizations can introduce systematic usage monitoring and cost-aware scheduling mechanisms that regulate how and when AI resources are consumed. By tracking model usage, limiting unnecessary invocations, and prioritizing high-impact operations, such mechanisms help prevent uncontrolled resource growth and ensure that the economic benefits of automation are not offset by hidden operational costs as the system scales.
4.1.10. Complexity of Non-Deterministic and Environmental Faults
While the proposed framework addresses the generation of patches for deterministic software defects, it is crucial to acknowledge the challenges posed by faults that defy straightforward code correction. A significant class of failures falls under the category of Mandelbugs, faults characterized by complex activation mechanisms, involving system-internal environmental interactions or propagation delays, resulting in non-deterministic behavior that is difficult to isolate or reproduce (Grottke and Trivedi, 2005). Unlike Bohrbugs, which manifest consistently under a defined set of inputs, Mandelbugs often arise from transient conditions such as race conditions, thread scheduling, or specific timing lags, rendering standard reproduction scripts ineffective.
Furthermore, a substantial subset of reported failures does not stem from incorrect source code logic but from environment-dependent and stateful conditions. These include transient issues such as network latency or service timeouts that resolve without intervention, configuration dependencies where bugs manifest only under specific system or database states, and resource constraints related to hardware limitations or memory exhaustion.
In a real example taken from (Trivedi et al., 2011), an aging-related Mandelbug causes non-deterministic server crashes in a tax system due to the cumulative exhaustion of the JVM heap during large-scale data processing. The failure probability increases over time as the system state degrades, requiring manual interventions like heap size adjustments and periodic restarts to restore stability.
In another example (Trivedi et al., 2011), timing-related Mandelbug occurs in a stock exchange system when a modification request is erroneously rejected because its corresponding order entry is still pending in a secondary processing queue. The failure is non-deterministic and load-dependent, manifesting only when specific timing conditions cause the order book update to lag behind incoming modification attempts.
Severe disk exhaustion can lead to physical hardware limitations where the existing storage infrastructure is no longer sufficient to handle the data load. These scenarios require manual physical maintenance, such as the physical installation of new drives or the removal of legacy hardware, to provide the necessary capacity for the system to resume operations.
Addressing these bugs requires a paradigm shift from static code analysis to dynamic system observability. Since the root cause often lies outside the codebase (e.g., in the deployment environment or infrastructure), automated patch generation techniques focused solely on syntax and logic modification are insufficient. Consequently, handling these failures requires not only code repair mechanisms but also operational interventions to address infrastructure and resource constraints. Therefore, the mitigation of Mandelbugs and environmental faults is considered outside the scope of this study, which focuses on the automated repair of deterministic logic defects.
4.1.11. Organizational Resistance, Role Displacement, and Change-Management Risk
A key socio-technical risk of the proposed workflow is organizational resistance driven by perceived role displacement. As summarized in Table 2, several stakeholders’ responsibilities shift from hands-on execution (e.g., manual triage, reproduction, test execution, and operational deployment activities) toward supervision and exception handling in a human-in-the-loop (HIL) setting. While this redesign is intended to reduce TTR and coordination overhead, it can trigger practitioner backlash, fear of job loss, and concerns about deskilling—particularly for roles where automation is explicitly positioned as taking over core activities (e.g., customer support’s validation/reproduction and no-code fixes; testers’ execution-heavy verification tasks; ops’ deployment pipeline work).
This risk is not only a cultural adoption issue; it can directly impact system performance. If practitioners distrust automated outputs or perceive the system as replacing rather than augmenting them, they may (i) resist using the workflow, (ii) bypass automation (reverting to “shadow processes” outside the tracker), (iii) provide minimal feedback needed for continuous improvement, or (iv) over-correct by adding excessive manual checkpoints—negating the intended efficiency gains. This concern complements the broader set of challenges already discussed under “Future Challenges and Risks,” which currently emphasize technical and governance limitations (e.g., cascading errors, accountability, evaluation, privacy, and cost).
Mitigating this challenge requires explicit change-management and job-redesign strategies alongside technical design. Concretely, future implementations should: (1) define “automation boundaries” per role (what is automated vs. what remains human-owned) and communicate them transparently; (2) frame the new responsibilities as capability uplift (e.g., “test reviewer,” “support supervisor,” “ops governor”) and invest in reskilling paths aligned with those responsibilities; (3) adopt phased rollouts and pilots where automation starts as decision support before becoming default, with measurable opt-out criteria; and (4) ensure the human-in-the-loop (HIL) checkpoints make people’s work easier, not harder, by clearly explaining why the system made a recommendation and by specifying when a person should step in or escalate the issue, so responsibility stays clear and manageable.
@@ 4.2. Implications for Practitioners
The proposed framework is not only a technical contribution but also a practical blueprint for how modern software teams can restructure their bug management processes in the presence of intelligent automation. Its successful adoption depends on organizational context, team structure, existing development practices, and the degree to which human oversight is integrated into automated workflows. The following implications highlight key considerations for practitioners, focusing on how the system can be adapted, customized, and integrated into real-world development environments to maximize its impact on productivity, reliability, and software quality.
@@ 4.2.1. Adapting to Diverse Project Structures
Our proposed bug tracking system assumes a certain team structure with dedicated roles for PMs, developers, and other team members. However, we acknowledge that this may not align with every project. The organizational structure of software development teams can vary significantly across different projects and companies. It is important for practitioners to adapt our system to fit their specific needs.
@@ 4.2.3. Modular System Architecture and Process Optimization
The components of our bug tracking system are designed with a modular architecture, meaning each part, from bug report enhancement and reproduction to fault localization and patch generation, can be run either independently or as part of a single, unified workflow. This is a critical implication for practitioners, as it allows for significant flexibility and process optimization based on a project’s unique needs and constraints.
@@ 4.2.4. System Integration and Toolchain Adapters
Successfully integrating such a bug tracking toolchain into existing systems requires a strategic approach focused on building flexible adapters and connectors. These components serve as a crucial layer, allowing the system to communicate with and leverage existing development tools without requiring a complete overhaul of an organization’s infrastructure. It is essential to develop these connectors for critical parts of the toolchain, such as VCSs like Git, which would allow the system to pull code and submit AI-generated patches as pull requests. Similarly, adapters for integrated development environments like Visual Studio Code [^38] https://code.visualstudio.com/ or IntelliJ [^39] https://www.jetbrains.com/idea/ would provide developers with real-time AI recommendations and debugging insights directly within their coding environment. Furthermore, building connectors for CI/CD platforms such as Jenkins [^40] https://www.jenkins.io/ or GitHub Actions is vital for automating the testing and validation of AI-generated fixes, ensuring that only verified code is merged. Finally, adapters for pre-existing bug tracking systems like Jira would allow for the seamless import of bug reports, enabling a phased and non-disruptive adoption. This approach ensures that the new system acts as a powerful enhancement to the existing ecosystem, maximizing the return on the integration effort.
@@ 4.3.3. Agent Capabilities as Research Problems
As illustrated in Figure 2, each capability represented in the workflow, such as bug report enhancement, reproduction, classification, localization, patch generation, and verification, can itself be viewed as a distinct research problem. Advances in NLP can refine bug report enhancement, program analysis techniques can strengthen localization, and test generation research can improve reproduction and verification. Researchers can therefore explore each of the boxes in Figure 2 as independent research topics, while also examining how these functions interoperate within an end-to-end framework that spans the entire bug tracking system.
@@ 4.3.4. Domain-Specific Challenges
The effectiveness of automated bug tracking will vary across domains. For example, in on-demand delivery systems, bugs often occur immediately, making them straightforward to detect and reproduce. In contrast, interactive environments such as gaming may require long execution traces or rare conditions before a bug surfaces (Yapağcı et al., 2025), complicating both reproduction and verification. These contrasts suggest that domain-specific adaptations of automation strategies are a promising research direction.
@@ 5. Conclusion
6. Acknowledgements
This work was supported by TÜBİTAK (The Scientific and Technological Research Council of Turkey) under the 1001 Scientific and Technological Research Projects Funding Program, Project No. 125E371. The authors gratefully acknowledge TÜBİTAK for its support.